OpenStudy (anonymous):

let f(x)=5^x use a small interval to estimate f^1(2) how do i do this

OpenStudy (anonymous):

\[f(x)=5^x\] and you want \[f'(2)\] is my guess

OpenStudy (anonymous):

right f prime of 2-i don't know how to type that either-but it looks like you do!

OpenStudy (anonymous):

then i think what you are being asked to do is compute \[\frac{5^{2+h}-5^2}{h}\] for small values of h, so crank up the calculator and try \[\frac{5^{2.1}-25}{.1}\] and \[\frac{5^{2.01}-25}{.01}\] etc

OpenStudy (anonymous):

if you want to cheat and go to the "exact" answer i can tell you that is it \[25\ln(5)\]

OpenStudy (anonymous):

cheers thanks.

OpenStudy (anonymous):

cheers thanks.

OpenStudy (anonymous):

thanks but i don't have to go that far yet...that will probably be in tomorrow's class! big help though-much appreciated!

OpenStudy (anonymous):

yw